### v1

# : An Open-Source SRAM Yield Analysis and Optimization Benchmark Suite

Abstract
Static Random-Access Memory (SRAM) yield analysis is essential for semiconductor innovation, yet research progress faces a critical challenge: the significant disconnect between simplified academic models and complex industrial realities. The absence of open, realistic benchmarks has created a reproducibility crisis, where promising academic techniques often fail to translate to industrial practice. We present OpenYield, a comprehensive open-source ecosystem designed to address this critical gap through three core contributions: (1) A realistic SRAM circuit generator that uniquely incorporates critical second-order-effect parasitics, inter-cell leakage coupling, and peripheral circuit variations, which are typically omitted in academic studies but decisive in industrial designs. (2) A standardized evaluation platform with a simple interface and implemented baseline yield analysis algorithms, enabling fair comparisons and reproducible research. (3) A standardized SRAM optimization platform, demonstrating OpenYield’s utility in enhancing SRAM design robustness and efficiency, providing a comprehensive benchmark for optimization algorithms. OpenYield creates a foundation for meaningful academia-industry collaboration, accelerating innovation in memory design. The framework is publicly available on OpenYield:URL.
I Introduction
Static Random-Access Memory (SRAM) has emerged as the dominant on-chip memory technology in modern integrated circuits, consuming up to 70% of die area in advanced AI accelerators and 60% in typical mobile System-on-Chips (SoCs) [1]. As semiconductor manufacturing pushes into sub-5nm nodes, the economic stakes of SRAM yield have become paramount—a mere 1% yield loss in high-volume production translates to millions in revenue impact, making accurate yield prediction not just a technical challenge but a critical business imperative for the $580 billion semiconductor industry [2].
The computational challenge of SRAM yield analysis is daunting. Modern memory arrays require failure rates below to be viable, yet directly observing such rare events would require over Monte Carlo simulations—computationally infeasible even with today’s massive computing resources [3]. This fundamental limitation has spurred decades of research into more efficient techniques, from variance reduction methods like importance sampling [4] and statistical blockade [5], to modern machine learning approaches [6, 7] and surrogate modeling techniques [8, 9]. Each promises to dramatically reduce simulation costs while maintaining accuracy.
Yet despite this rich algorithmic innovation, a troubling pattern has emerged: techniques that show exceptional promise in academic publications routinely fail when applied to industrial designs. Our comprehensive review of 45 recent SRAM yield analysis papers reveals the root cause—over 85% rely on dramatically oversimplified models that omit the very physical effects in real circuits. This systematic disconnect between academic models and industrial reality has created what we term a “reproducibility crisis” in SRAM yield analysis.
The gap between academic models and industrial reality is qualitative, not merely quantitative. Real SRAM arrays exhibit complex phenomena that fundamentally alter behavior: parasitic resistances slow bitline discharge by 4-6 [10]; leakage from hundreds of unselected cells degrades read margins by 12% [11]; sense amplifier offset variations of 15-30mV shift failure boundaries unpredictably [12]; and position-dependent IR drops create systematic variations across arrays [13]. When combined, these effects create failure modes that simply cannot exist in simplified models. This modeling gap has severe practical consequences. The semiconductor industry spends $2-3 billion annually on design respins, with yield issues driving 30% of these costs. More critically, the disconnect undermines research itself: algorithms showing similar performance on simplified benchmarks may differ by orders of magnitude on real circuits; optimization techniques that appear optimal in studies fail catastrophically in practice; and without standardized, realistic benchmarks, fair comparison of approaches becomes impossible.
To address this systemic challenge, we present OpenYield: an open-source ecosystem for realistic SRAM yield analysis and optimization. Our vision is to democratize access to industrial-quality SRAM modeling, enabling the academic community to develop and validate techniques on circuits that actually represent the challenges faced in modern semiconductor design. OpenYield makes three core novelties:
A Python-based, scalable SRAM generator that configures any-sized array with high-quality modeling of parasitics of networks, inter-cell leakage, and peripheral circuit variations—critical effects absent in academic studies. Validation shows models without parasitics underestimate read delay and power by 18.7 and 16.7; models without peripherals underrate by 4.8 and 5.4.
A Standardized Yield Analysis Platform with reference implementations of baseline algorithms (Monte Carlo, Importance Sampling variants, ML approaches) operating on realistic circuits, enabling reproducible benchmarking across different SRAM configurations.
A Standardized SRAM Optimization Platform with five state-of-the-art optimization algorithms for transistor sizing under realistic conditions, enabling fair comparison and demonstrating up to 65% SNM improvements and 15% area reduction over baseline designs.
Like ImageNet [14] for computer vision or SPEC [15] for computer architecture, OpenYield establishes standardized benchmarks that enable reproducible research and accelerate innovation in memory design. The complete framework is publicly available on OpenYield:URL.
II Background
II-A Physical Origins of the SRAM Yield Challenge
Process variations in semiconductor manufacturing create the fundamental yield challenge in SRAM design. As transistors shrink to atomic scales, random dopant fluctuation becomes inevitable—a 20nm transistor contains only 100 dopant atoms in its channel, making each atom’s position statistically significant [16]. These atomic-level variations manifest as threshold voltage () mismatch between transistors, with standard deviations following Pelgrom’s law: , where ranges from 2-4 mV·µm in modern processes [17].
SRAM cells amplify these variations through their ratioed design. Read stability requires the pull-down transistor (PD) to overpower the access (PG) transistor, preventing the stored node from flipping. Write ability demands the opposite—the PG transistor must overpower the pull-up (PU). These conflicting requirements create narrow design margins that variations easily violate. A 30mV shift well within 3 can push a marginal cell into failure. The statistics compound mercilessly. A 1Mb array contains over 6 million transistors. Even with 6 process control, several thousand transistors will exhibit extreme variations. Since a single failing cell renders the entire array unusable, achieving acceptable yield requires individual cell failure rates below , deep in the statistical tail, where rare events dominate.
II-B Formulation of Yield Analysis and SRAM Optimization
Yield analysis quantifies the probability that a manufactured array functions correctly despite process variations. For an array with identical and independent cells, where each cell has a failure probability of , the overall yield can be expressed as . For small failure probabilities, this can be approximated as . This simple relationship shows how even a small per-cell failure probability can dramatically impact the yield of large arrays.
This exponential relationship drives the stringent requirements. Achieving 90% yield for a 1Mb array demands . Monte Carlo simulation provides the reference method for estimating . Generate samples from the process distribution, simulate each, calculate failures as .For , this demands simulations.
SRAM optimization navigates competing objectives under yield constraints, formulated as
where represents transistors’ widths (W) and lengths (L) in a SRAM cell, are objectives (e.g., power, area, noise margin), and can be yield-related constraints. Critically, these relationships change dramatically based on which physical effects the model includes.
II-C The Modeling Gap: From Simplification to Crisis
Academic SRAM research suffers from systematic oversimplification. Real arrays exhibit failure modes that simplified models cannot capture:
Distributed parasitics create position-dependent behavior. A 256-row bitline presents fF capacitance and resistance, not as lumped elements but distributed along the wire. Cells at the array’s far end see 4-6 slower discharge rates. Simple RC models miss the incomplete swing and non-monotonic delay that distributed effects create.
Inter-cell coupling introduces data-pattern dependencies. With 255 unselected cells per bitline, cumulative leakage reaches microamps—comparable to the selected cell’s read current at low voltages [11]. The resulting margin degradation depends on stored data patterns, creating failures invisible to single-cell analysis.
Peripheral circuits often dominate yield loss. Sense amplifier (SA) offset voltage ( 20mV) directly subtracts from read margin [12]. Wordline driver skew creates systematic timing variations across columns. These effects require modeling the complete SRAM array, not just the core cell.
Yield analysis algorithms validated on simplified circuits may produce order-of-magnitude errors on real arrays. Without intervention, this gap threatens the entire research enterprise. Algorithms cannot advance without realistic validation. Industry cannot adopt techniques proven only on toy models. The field needs standardized, realistic benchmarks—exactly what OpenYield provides.
II-D SOTA SRAM Tools and Their Limitations
OpenRAM [18] compiles complete SRAMs from specification to GDS with diverse configurations and manufacturable layouts, but lacks variation modeling and yield analysis infrastructure.
SRAM22 [19] provides automated embedded SRAM generation with compact layouts, but similarly offers no statistical analysis or yield optimization features.
AnalogGym [20] benchmarks general analog synthesis algorithms but cannot capture SRAM-specific phenomena like massive parallelism, statistical extremes, and critical layout dependencies.
Commercial tools accurately model these effects but remain inaccessible due to proprietary licenses, black-box implementations, and high costs that exclude academic research. OpenYield addresses this gap by combining industrial-quality models with open-source accessibility, enabling reproducible research and advancing SRAM yield analysis.
III
: Design and Implementation
The OpenYield ecosystem follows a modular architecture depicted in Fig. missing 1. The core SRAM Circuit Generator produces PySpice-based netlists [21] that incorporate user configurations and realistic physical effects. These netlists are analyzed by an engine using open-source SPICE tools like Xyce [22].
Two principal blocks leverage the simulation insights: Yield Analysis Algorithms (Sect. missing III-B) implement statistical techniques for yield estimation and failure probability assessment, while the Design Optimization Framework (Sect. missing III-C) automatically adjusts design parameters (e.g., bitcell transistor sizes) to enhance SRAM robustness and performance.
Users interact with OpenYield by providing comprehensive inputs including SRAM configurations (arbitrary array dimensions, memory cell types from 6T standard with 8T/10T provisions), technology characteristics (currently FreePDK45 [23] support), design parameters, process variation statistics (local intra-die random variations applied independently to each transistor using Gaussian distributions), and optimization goals for Monte Carlo simulations. The platform delivers multiple outputs: SPICE netlists (.sp format) ensuring broad simulator compatibility, detailed yield reports, failure probability estimates, characterized performance metrics, and optimized design parameters. Monte Carlo parameterization can be embedded directly in model cards or exported as data tables for importance sampling-based yield analysis algorithms. The platform follows open-source principles, encouraging community contributions and extensions.
III-A Hierarchical SRAM Generator
The cornerstone of OpenYield is its novel SRAM circuit generator, engineered to produce unified benchmark circuits that transcend common academic simplifications and approach industrial SRAM complexities.
III-A1 Circuit Architecture
As illustrated by Fig. missing 2, the 6T SRAM cell comprises cross-coupled inverters (M0-M1, M4-M5) for storage and PG (M2-M3) for read/write control. Process variations, particularly random dopant fluctuation, create threshold voltage mismatches that drive yield loss through specific failure mechanisms:
Read Upsets occur when bitline BLB discharge through M3-M1 causes node V(QB) of M4-M0 inverter, destructively flipping the cell state. Read Static Noise Margin (RSNM) quantifies the cell’s immunity to this—insufficient RSNM (typically 50mV) directly causes the failure.
Write failures happen when M2 cannot pull node Q below the M5-M1 trip point against M4’s pull-up action. Write Static Noise Margin (WSNM) measures the cell’s ability to flip during write operations—inadequate WSNM (50mV) prevents successful data writing.
Hold failures occur when the cell loses stored data due to leakage currents or VDD droops in standby mode, when both bitlines float and the wordline is inactive. Hold Static Noise Margin (HSNM) characterizes data retention stability—insufficient HSNM (100mV) leads to spontaneous data corruption during idle periods.
Access time failures result when the bitline differential cannot overcome the within the allocated timing window. Both read delay (, including time to develop sufficient bitline differential) and write delay (, including time to flip the cell state) must meet timing constraints—violations cause functional failures regardless of noise margins.
All these failure mechanisms can be simulated and analyzed through OpenYield. The generator constructs complete array hierarchies including: core cells in configurable row/column arrangements, wordline drivers sized for distributed loads, column multiplexers with shared sense amplifiers, precharge/write driver circuits, sense amplifiers with realistic offset distributions, and configurable control logic. This comprehensive approach distinguishes OpenYield from existing academic generators that produce idealized circuits.
III-A2 Second-Order Physical Effects Modeling
OpenYield integrates critical physical effects that profoundly influence yield in advanced technologies:
Parasitic Network Generation: Distributed -shaped RC segments model vital conductive paths (bitlines, wordlines) rather than oversimplified lumped capacitances. Users can specify resistance/capacitance values for each subcircuit, with defaults based on typical SRAM layout topologies.
Inter-Cell Leakage Coupling: Models various leakage mechanisms (subthreshold, gate leakage) to simulate neighboring cell impacts on actively accessed cells—particularly significant under low VDD conditions.
Peripheral Circuit Variation: Instantiates realistic models for sense amplifiers (with systematic/random offset voltage distributions), wordline drivers (with delay variations and skews), and precharge/write drivers (with voltage level and timing variations). These peripheral variations are modeled conjunctively with core cell variations for holistic yield assessment.
III-B Yield Analysis Platform
OpenYield provides optimized implementations of established yield analysis methods, creating standardized benchmarks for algorithm comparison and reproducible research.
III-B1 Monte Carlo Reference
Standard Monte Carlo serves as the accuracy baseline, automating variation sampling, SPICE simulation, and pass/fail evaluation. Enhanced with stratified sampling and control variates, the implementation reduces required samples by - while maintaining unbiased estimation—essential for validating advanced methods.
III-B2 Importance Sampling Variants
Four IS methods address different yield analysis challenges. Mean-shifted IS (MNIS) [4] translates distributions toward the most probable failure point: subject to failure boundary . This achieves speedup when single modes dominate, particularly for read failures driven by sense amplifier mismatch. Adaptive Compressed Sampling (ACS) [24] builds sparse failure representations using compressed sensing, reducing samples by 90% for smooth failure boundaries. Adaptive IS (AIS) [25] iteratively refines proposal distributions through cross-entropy minimization, automatically discovering multiple failure modes. High-dimensional Sparse Compressed Sampling (HSCS) [26] exploits sparsity to scale efficiently to hundreds of parameters, crucial for full-array analysis with systematic variations.
III-C Standardized SRAM Optimization Platform
OpenYield’s optimization platform provides a standardized benchmarking environment for SRAM transistor sizing algorithms, enabling fair comparison of different optimization approaches under identical realistic circuit conditions. The platform supports flexible optimization formulations including single-objective, multi-objective, and constrained problems, allowing users to combine different objectives and constraints based on their specific requirements.
III-C1 Flexible Optimization Formulation
The platform accommodates diverse optimization scenarios through configurable objective functions and constraints. For demonstration, we present a single-objective formulation using a 6T SRAM cell within a array configuration. The optimization explores design parameters : device types with different threshold voltages (vtl/vtg/vth), gate widths (W) of M0/M2/M4, and gate lengths (L). Parameters W and L range from 0.5 to 1.5 nominal values with 5nm steps. The optimization problem can be formulated as:
where
Users can easily customize objectives (power, area, noise margins), constraints (timing, yield requirements), and combine multiple formulations to address specific design challenges under realistic circuit conditions.
III-C2 Integrated Optimization Algorithms
The platform implements five state-of-the-art optimization algorithms for comprehensive benchmarking: Constrained Bayesian Optimization (CBO) [27] using Gaussian Process surrogates with Expected Hypervolume Improvement acquisition; Particle Swarm Optimization (PSO) [28] with adaptive parameters for population-based search; Simulated Annealing (SA) [29] with exponential cooling schedules; Sequential Model-based Algorithm Configuration (SMAC) [30] for automated hyperparameter optimization; and RoSE-Opt [31] combining Bayesian optimization with reinforcement learning for adaptive exploration.
Each algorithm operates under identical evaluation budgets and circuit conditions, enabling fair performance comparison across different optimization paradigms. The platform provides a unified evaluation framework where all algorithms utilize OpenYield’s realistic circuit models. Each optimization run performs an identical number of SPICE simulations, ensuring fair computational comparison. Comprehensive metrics track convergence characteristics, Pareto front quality, and final design performance. The platform’s integration with OpenYield’s realistic models ensures optimized designs account for the complex failure mechanisms dominating manufactured arrays, while enabling reproducible comparison of optimization algorithm effectiveness.
IV Experimental Results
All SRAM circuits for this evaluation were generated using the OpenYield framework. To ensure reproducibility, we primarily targeted a generic FreePDK45nm Technology Model. The nominal supply V. The default design parameters (all Ws and Ls refer to that in OpenRAM [18]). Process parameters and their variations, including threshold voltage without bias (vth0), low-field mobility (µ0), and offset voltage in subthreshold region (voff), were chosen based on typical values with . Computationally intensive SPICE simulations were executed through a parallel version of Xyce[22] to ensure timely completion.
Our primary experimental metrics are explained as follows. Yield, , where is the failure probability from unmet operational criteria (e.g., timing metrics). Static Noise Margins (SNM), including RSNM, WSNM, and HSNM, characterized via standard methods (e.g., largest square embedded within the lobes of butterfly curves) [32, 2]. 3(a) presents the HSNM curve. The outputs of the left- and right-side of inverters are denoted as V(V1) and V(V2) in Fig. missing 3. We rotate all SNM curves by 45∘ to explicitly depict the largest noise margin. A typical RSNM butterfly curve has a similar shape to HSNM (waveform is omitted). Note that the RSNM value is always smaller than HSNM due to BL or BLB cannot be ideally high or low due to the turned-on access transistors. 3(b) illustrates the write margin. The asymmetry of WSNM is caused by the write drivers. More space between the 2 curves makes it easier to write this cell.
Read Time (), from wordline (WL) assertion to a sufficient bitline (BL/BLB) differential for sense amplifier reliability. 4(a) depicts a typical read cycle of a 324 SRAM array with full peripheral circuits. Initially, V(BL3) and V(BLB3) are precharged high. Upon assertion of V(WL31), the cell begins to discharge BL3 due to V(CELL:Q3). The time taken to develop a differential voltage is a key component in . After a sufficient voltage differential is established (e.g., 250mV), SAE is assertion to activate the sense amplifier, converting the analog voltage swing to a logical level. Write Time (), from WL assertion and active BL/BLB drive until internal storage nodes successfully flip (V(CELL:Q)). A representative write cycle is shown in 4(b). After BL3 and BLB3 are driven to complementary target values, the write operation is initiated by asserting Write Enable (WE) and pulling up WL31. A transition state of CELL:Q/QB kicks in, and finally, they reach a new stable state. Typically, is much shorter than in a balanced SRAM design, this is because (1) shorter interconnection of CELL:Q/QB has smaller parasitics than that of the read port (i.e., BLs); (2) the PU MOS (M4/M5) is designed to be weaker to increase the write stability.
Other performance metrics include Power Consumption, covering static and dynamic power, and Cell Area, calculated based on the DRC rules in FreePDK45.
IV-A Experiment 1: Scalability of the OpenYield Generator
This subsection demonstrates the versatility of the OpenYield generator by showcasing its ability to produce SRAM arrays of varying sizes and complexities.
Fig. missing 5 illustrates the impact of scaling the number of rows on key performance metrics for a 4-column SRAM array. Specifically, it shows the mean access delay and average access power as the number of rows varies from 8 to 256. The data indicates a clear trend: increases from approximately 0.14ns for an 8-row array to 3.27ns for a 256-row array. Similarly, the average during read operations scales from 14.1µW to 434.5µW across the same range of row sizes. In contrast, shows minimal variation with the number of rows. , while also scaling with array size (2.7µW 8.4µW), is substantially lower than read power.
IV-B Experiment 2: Quantifying Second-Order Effects
To quantify the impact of a key second-order effect, we performed comparative MC simulations on SRAM arrays of varying row sizes. One set of simulations included detailed parasitic resistance and capacitance (RC) models for interconnects, while the other intentionally omitted them, representing an idealized scenario. We focused these comparisons on read operations, as read delay and stability are often the most critical performance limiters in SRAM design, as mentioned before.
Impact of Interconnect Parasitics: The inclusion of parasitic RC elements associated with the interconnects is expected to substantially affect both access delay and power consumption, particularly as array dimensions increase.
The results presented in Fig. missing 6 clearly demonstrate the profound impact of interconnect parasitics. For instance, for a 64-row array, the mean read access delay increases from approximately 0.067ns without parasitics to 0.843ns when parasitics are included—an increase by a factor of roughly 12.6. This discrepancy becomes even more pronounced for larger arrays; for a 256-row array, the delay with parasitics is approximately 18.7 greater than without (3.273ns 0.175ns).
Similarly, average is significantly underestimated when parasitics are ignored. For the 64-row array, including parasitics increases power by (6.87µW 109.47µW). For the 256-row array, rises by ( 25.99µW 434.51µW).
Impact of Inter-cell Leakage (Leakage Noise): To characterize the impact of this leakage noise, we simulated a 644 SRAM array across a range of supply voltages (VDD from 0.45V to 1.0V). We investigated the read delay under two scenarios representing different leakage conditions from idle cells in the same column: (1) all idle cells storing ’0’, and (2) all idle cells storing ’1’.
Our first finding is that the read access delay is strongly dependent on the supply voltage, decreasing from over 6ns at V to approximately 0.84ns at V for both idle-cell data patterns. At lower supply voltages, the effect of leakage noise is more pronounced. For example, at V, the read delay when idle cells store ‘0’ (6.70ns) is about 0.32ns (or ) longer than when idle cells store ‘1’ (6.38ns). However, this difference diminishes as VDD increases, becoming negligible at V and above.
Impact of Peripheral Circuit Variations: The variations of peripheral circuits play a crucial role in overall SRAM performance and reliability. To demonstrate their impact, we compare simulations performed with a simplified model versus the model with full peripherals for a 6432 SRAM array.
Tab. missing I summarizes key delay and power metrics for read and write operations under these two modeling conditions.
| Operation | Type | Metric | w/o Peri. | w/ Peri. |
| Read | Delay (ns) | N/A | 1.0853 | |
| N/A | 0.2283 | |||
| 0.8523 | 4.0864 | |||
| Power (mW) | avg | 0.8757 | 4.7073 | |
| dyn | 3.4760 | 6.9954 | ||
| static | 0.0094 | 0.0127 | ||
| Write | Delay (ns) | N/A | 0.7155 | |
| N/A | 0.2319 | |||
| 0.0974 | 0.3799 | |||
| Power (mW) | avg | 0.0254 | 0.1577 | |
| dyn | 0.0525 | 0.3546 | ||
| static | 0.0057 | 0.1907 |
For read operations, critical delay components associated with peripherals like precharge time (ns) and wordline drive time (ns) are inherently captured. Furthermore, the bitline related delay () shows an increase of with peripherals (0.8523ns 4.0864ns). This highlights the considerable loading and delay contributions from components like sense amplifiers and column multiplexers. Read power consumption () is also markedly higher with peripherals: average increases by (0.8757mW 4.7073mW), and dynamic by .
For write operations, similar trends are observed. Delays specific to write drivers (ns) and wordline drivers (ns) are accounted for with the full model. The cell’s internal node switching time () also has a increase (0.0974ns 0.3799ns) when peripheral interactions are considered. Write power () metrics show even more dramatic increases: average is higher, dynamic is higher, and notably, static sees an increase of over when peripheral circuits are included. These comparisons underscore that omitting detailed models of peripheral circuits and their variations can lead to highly optimistic and inaccurate estimations of SRAM performance, particularly for delay and power.
IV-C Experiment 3: Validation of Baseline Yield Algorithms
This experiment validates the baseline yield analysis algorithms in OpenYield—Standard Monte Carlo (MC) and Importance Sampling (IS)—and demonstrates OpenYield as a platform for fair comparison of analysis techniques on realistic circuits. The test circuits include: (1) a single 6T cell with 18 process variation parameters; (2) a 32 array with 108 parameters; and (3) a 322 array with 1152 parameters. Write delay serves as the performance metric across all experiments. The figure of merit is , where represents the standard deviation of the estimated failure rate, serving as the termination criterion.
| Method | Cell | Array | Array | |||||||||
| Fail | FoM | #Sim | Time(s) | Fail | FoM | #Sim | Time(s) | Fail | FoM | #Sim | Time(s) | |
| MC | 2.4e-4 | 0.100 | 102,400 | 7,567 | 2.0e-4 | 0.070 | 130,000 | 24,492 | 2.2e-4 | 0.076 | 128,500 | 199,447 |
| MNIS[4] | 2.1e-4 | 0.017 | 12,450 | 1,075 | 2.7e-3 | 0.049 | 8,600 | 1,771 | 9.6e-3 | 0.049 | 5,540 | 10,178 |
| AIS[25] | 6.1e-4 | 0.056 | 10,050 | 934 | 9.3e-4 | 0.046 | 9,200 | 1,442 | 1.6e-3 | 0.069 | 6,700 | 7,091 |
| ACS[24] | 1.7e-3 | 0.018 | 5,600 | 946 | 3.8e-5 | 0.049 | 4,600 | 1,035 | 6.2e-5 | 0.049 | 5,800 | 10,050 |
| HSCS[26] | 5.1e-2 | 0.049 | 11,900 | 1,469 | 5.5e-4 | 0.049 | 17,900 | 1,153 | 1.6e-2 | 0.049 | 11,300 | 10,312 |
Results for the single 6T cell are shown in Tab. missing II. MNIS [4] achieves near-MC accuracy with less than 10% error while delivering 7 speedup. AIS [25] provides a moderate balance between precision and efficiency, though with higher error than MNIS. For the 32 case (Fig. missing 7 and Tab. missing II), ACS [24] uniquely surpasses MC accuracy with sub-10% error while achieving 23.7 speedup. However, MNIS exhibits significantly degraded performance with error rates 10 higher than MC, revealing limited generalizability. AIS and HSCS show moderate precision but require substantially more simulations. In the largest case, MC requires over 128,000 simulations and 55 hours of runtime. ACS achieves comparable accuracy with only 5,800 simulations—a 19.8 speedup while maintaining stable performance. HSCS achieves the fastest runtime but suffers a drastic accuracy loss, with failure estimates over an order of magnitude higher than MC. MNIS and AIS also display unstable results.
These results establish OpenYield as a transparent environment for benchmarking yield analysis algorithms, enabling reliable comparisons within the research community.
IV-D Experiment 4: Design Improvement via Optimization
In this sub-section, we focus on multi-objective transistor sizing optimization to improve cell robustness while maintaining performance constraints. We implement a control study using simplified models (neglecting second-order effects) and re-evaluate the “optimal” design in [18]. The optimization objective can be found in Eq. (missing) 2.
We evaluate five state-of-the-art optimization algorithms, each performing 400 SPICE simulations: (1) CBO [27] using Gaussian Process surrogates with Expected Hypervolume Improvement acquisition and constraint handling; (2) PSO [28] with population size 20, adaptive inertia weight (), and cognitive/social parameters (); (3) SA [29] with exponential cooling schedule (, , ); (4) SMAC employing sequential model-based configuration; and (5) RoSE-Opt combining Bayesian optimization with reinforcement learning using Proximal Policy Optimization agents for adaptive parameter exploration.
Fig. missing 10 shows the Pareto fronts and Fig. missing 10 compares convergence characteristics across all algorithms. Tab. missing III shows the OpenYield-optimized design (RoSE-Opt) achieves significant improvements: 65% enhancement in read SNM (0.17V 0.28V), 39% improvement in write SNM (0.79V 1.10V), and 15% area reduction (0.61µm2 0.52µm2) while maintaining timing constraints. The composite FoM increases from 10.20 to 10.43, demonstrating balanced multi-objective improvement. SMAC and CBO show competitive performance with different exploration-exploitation trade-offs, while PSO and SA exhibit more stochastic convergence patterns.
Critically, designs optimized using simplified models exhibit poor performance when evaluated under realistic conditions, with some violating timing constraints or showing degraded stability margins. This validates the necessity of high-fidelity modeling during optimization, positioning OpenYield as an essential tool for robust SRAM design.
| Parameter/Metric | w/o Opt. | w/ RoSE-Opt | |
| Design Params | NMOS Model | NMOS_VTG | NMOS_VTH |
| Type | PMOS_VTG | PMOS_VTG | |
| W_PD (µm) | 0.205 | 0.170 | |
| W_PU (µm) | 0.090 | 0.124 | |
| W_PG (µm) | 0.135 | 0.075 | |
| L (nm) | 50 | 74 | |
| Performance Metrics | (0.5ns) | 0.4533 | 0.4374 |
| (0.5ns) | 0.0903 | 0.0583 | |
| (µW) | 13.74 | 14.37 | |
| (µW) | 0.86 | 1.14 | |
| HSNM (V) | 0.33 | 0.33 | |
| RSNM (V) | 0.17 | 0.28 | |
| WSNM (V) | 0.79 | 1.10 | |
| Area (µm2) | 0.61 | 0.52 | |
| FoM | 10.20 | 10.43 | |
V Conclusion
OpenYield bridges the academia-industry gap in SRAM yield analysis through the first open-source platform with industrial-quality modeling. Though limited to FreePDK45 and 6T cells, it enables reproducible benchmarking and meaningful collaboration between researchers and industry. Future extensions include extensions to advanced nodes and layout-aware analysis and optimization.